OpenStudy (anonymous):

Give the exact and approximate solutions to three decimal places. y^2-8y+16=36 y=?

OpenStudy (anonymous):

Move the 36 over. y^2-8y-20=0 Factor (quadratic formula). The rest you can do.. I am simply too lazy.

OpenStudy (anonymous):

(y+2)(y-10)=0

OpenStudy (anonymous):

so y=-2 and y=10

OpenStudy (anonymous):

y+2=0...y=-2 y-10=0...y=10

OpenStudy (anonymous):

yeah thats it ..lol

OpenStudy (anonymous):

so y=-2, 10

OpenStudy (anonymous):

exact and approximate? who writes these?

OpenStudy (anonymous):

justa check on...y.2x10=20...-10y+2y=8y...yxy=y^2

OpenStudy (anonymous):

oops sory i mean check on....2(-10)=-20...-10y+2y=-8y...yy=y^2

OpenStudy (anonymous):

thnx